# The new observations about the parameter-dependent Schrödinger-Poisson system

### v1

@@ 1. Introduction and Main Results
Under assumptions ${\bf(V)}$ and ${\bf(f)}$. Thne, there is a constant $\lambda^{*}>0$ such that if $\lambda\geq\lambda^{*}$, system (1.1) admits a non-trivial weak solution $u$ in the following sense
@@
$\Delta u+V(x)u=\frac{1}{4\pi}\int_{\mathbb{R}^{3}}\frac{u^{2}(y)}{|x-y|}\text{d}yu\ \ ~\text{in}~\mathbb{R}^{3}.$
The paper is organized as follows. In Section 2, we propose a new modified problem. From this new modification method, it can be seen that there is a close relationship between the solution of this modified problem and the solution of the original problem. In Section 3, we prove that the modified functional has critical points, and when the parameter is sufficiently large, the critical point is exactly the critical point of the original problem. In Section 4, we prove that similar results in the case of sign-changing potentials. In Section 5, we give the proof of the asymptotical behavior of non-trival solutions.
@@ 3. The existence of parameter-dependent Schrödinger-Poisson system
In this section, the non-trivial solution of systme (1.1) is obtained and give the proof of Theorem 1. Firstly, the mountain pass theorem is applied to prove the existence of critical points for the modified functional $I_{\epsilon}$. Secondly, the truncation of $\tilde{F}$ is removed when $\epsilon>0$ is sufficiently small by Moser iteration. Finally, the mountain pass critical point of the modified functional $I_{\epsilon}$ is actually the solution to the original system (1.1).
@@
According to Lemma 1, it follows from the mountain pass theorem in Willem [25] that there is a sequence $\{u_{n}\}_{n=1}^{\infty}\subset H^{1}_{r}(\mathbb{R}^{3})$ such that
@@
It follows from [12, Lemma 1] that
@@
Note that $\phi_{u_{n}}\rightarrow\phi_{u}$ in $L^{6}(\mathbb{R}^{3})$ and $u_{n}\rightarrow u$ in $L^{12/5}{(\mathbb{R}^{3})}$. Then
@@
By the Lemma 2, there is a $u_{\epsilon}\in H^{1}_{r}(\mathbb{R}^{3})$ such that $I_{\epsilon}(u_{\epsilon})=c_{\epsilon}$ and $I_{\epsilon}^{\prime}(u_{\epsilon})=0$. Therefore, $u_{\epsilon}$ is the weak solution of system (2.3). The following lemma enables us to remove the truncation (2.1).
@@
For any $R>1$, there exists $\epsilon^{\prime}(R)>0$ such that for $0<\epsilon<\epsilon^{\prime}(R)$, let $u_{\epsilon}$ be the weak solution of system (2.3) obtained in Lemma 2. Then there exists a positive constant $m_{0}$ independent of $\epsilon$ and $R$ such that
@@
For any $0<\epsilon<\epsilon_{*}(R)$, let $u_{\epsilon}$ satisfy $I_{\epsilon}(u_{\epsilon})=c_{\epsilon}$ and $I_{\epsilon}^{\prime}(u_{\epsilon})=0$. Using the similar method in Lemma 2, it implies
@@
It is easy to say that $\{u_{\epsilon}\}$ is a bounded sequence. Therefore, there exists $M_{0}>0$ independent of $\epsilon$ such that for all $\epsilon\in(0,\epsilon_{*}(R))$, it holds
@@
(3.11) $\int_{\mathbb{R}^{3}}\left(\nabla u_{\epsilon}\nabla\varphi+V(x)u_{\epsilon}\varphi\right)\text{d}x=\int_{\mathbb{R}^{3}}\phi_{u_{\epsilon}}u_{\epsilon}\varphi\text{d}x-\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\widetilde{f}_{R}\left(\epsilon u_{\epsilon}\right)\varphi\text{d}x.$
@@
$\quad\left|\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x-\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\widetilde{f}_{R}\left(\epsilon u\right)u\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x\right|$
$\leq\left(\int_{\mathbb{R}^{3}}\phi_{u}^{6}\text{d}x\right)^{\frac{1}{6}}\left(\int_{\mathbb{R}^{3}}\left|u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\right|^{\frac{6}{5}}\text{d}x\right)^{\frac{5}{6}}+\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\left|\widetilde{f}_{R}\left(\epsilon u\right)u\right|\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x$
@@
$\leq CM^{2}_{0}\left(\int_{\mathbb{R}^{3}}\left|u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\right|^{\frac{6}{5}}\text{d}x\right)^{\frac{5}{6}}+\delta\int_{\mathbb{R}^{3}}u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x+C_{\delta,R}\int_{\mathbb{R}^{3}}\epsilon\left|u\right|^{3}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x.$
Then take $\delta=\frac{V_{0}}{2}$ and $\epsilon(R)=C_{\frac{V_{0}}{2},R}^{-1}$ for $0<\epsilon\leq\epsilon(R)$, one has
$\quad\left|\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x-\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\widetilde{f}_{R}\left(\epsilon u\right)u\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x\right|$
$\leq C_{1}\left(\int_{\mathbb{R}^{3}}\left|w_{k}\right|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}+\frac{V_{0}}{2}\int_{\mathbb{R}^{3}}w_{k}^{2}\text{d}x+\epsilon C_{\frac{V_{0}}{2},R}\int_{\mathbb{R}^{3}}w_{k}^{2}\left|u\right|\text{d}x$
$\leq C_{1}\left(\int_{\mathbb{R}^{3}}\left|w_{k}\right|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}+\epsilon C_{\frac{V_{0}}{2},R}\left(\int_{\mathbb{R}^{3}}u^{6}\text{d}x\right)^{\frac{1}{6}}\left(\int_{\mathbb{R}^{3}}|w_{k}|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}+\frac{V_{0}}{2}\int_{\mathbb{R}^{3}}w_{k}^{2}\text{d}x$
@@
The proof of Theorem 1. According to Lemmas 1 and 2, for $0<\epsilon<\epsilon_{*}$, there exists a nontrivial weak solution $u$ satisfies
@@
Furthermore, take $\epsilon^{*}=\min\{\frac{1}{m_{0}^{2}},\epsilon_{*}\}$. For $0<\epsilon<\epsilon^{*}$, it follows from (3.10) that
@@
[…]
@@ 4. Schrödinger-Poisson system with the sign-changing potential
The purpose of this section is to obtain a nontrivial solutions of system (1.1) with the sign-changing potential. In this case, the corresponding functional $I_{\epsilon}$ defined in (4.1) is different from (2.4). Before proving our results, we shall introduce the suitable function space where the critical points of the modified functional $I_{\epsilon}$ belong to.
Define the working space as follows:
@@ 4.1. Cerami condition
In order to prove Theorem 2, we firstly need to show that the functional $I_{\epsilon}$ satisfies the Cerami conditions. To this end, we rewrite $I_{\epsilon}$ in the following form
(4.1) $I_{\epsilon}(u)=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+\tilde{V}(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u)\text{d}x-\frac{m}{2}\int_{\mathbb{R}^{3}}u^{2}\text{d}x.$
@@
Let $\{u_{n}\}_{n=1}^{\infty}\subset H^{1}_{V}(\mathbb{R}^{3})$ be a Cerami sequence for the functional $I_{\epsilon}$. There exists a constant $\epsilon_{**}>0$ such that if $\epsilon\in(0,\epsilon_{**})$, then $u_{n}\rightarrow u\text{ in }H^{1}_{V}(\mathbb{R}^{3})$ with $u\in H^{1}_{V}(\mathbb{R}^{3})$.
@@
Let $\{u_{n}\}$ be a Cerami sequence of $I_{\epsilon}$, that is, for some $c\in\mathbb{R}$,
$I_{\epsilon}(u_{n})\to c~\text{and}~(1+\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})})I_{\epsilon}^{\prime}(u_{n})\to 0\ ~\text{as}~n\to+\infty.$

### v3

@@ 1. Introduction and Main Results
Under assumptions ${\bf(V)}$ and ${\bf(f)}$. Then, there is a constant $\lambda^{*}>0$ such that if $\lambda\geq\lambda^{*}$, system (1.1) admits a non-trivial weak solution $u$ in the following sense
@@
$\Delta u+V(x)u=\frac{1}{4\pi}\int_{\mathbb{R}^{3}}\frac{u^{2}(y)}{|x-y|}\text{d}y~u\ \ ~\text{in}~\mathbb{R}^{3}.$
The paper is organized as follows. In Section 2, we propose a new modified problem. From this new modification method, it can be seen that there is a close relationship between the solution of this modified problem and the solution of the original problem. In Section 3, we prove that the modified functional has critical points, and when the parameter is sufficiently large, the critical point is exactly the critical point of the original problem. In Section 4, we prove that similar results in the case of sign-changing potentials. In Section 5, we give the proof of the asymptotical behavior of non-trivial solutions.
@@ 3. The existence of parameter-dependent Schrödinger-Poisson system
In this section, the non-trivial solution of system (1.1) is obtained and give the proof of Theorem 1. Firstly, the mountain pass theorem is applied to prove the existence of critical points for the modified functional $I_{\epsilon}$. Secondly, the truncation of $\tilde{F}$ is removed when $\epsilon>0$ is sufficiently small by Moser iteration. Finally, the mountain pass critical point of the modified functional $I_{\epsilon}$ is actually the solution to the original system (1.1).
@@
According to Lemma 1, it follows from the mountain pass theorem in Ambrosetti and Rabinowitz [1] that there is a sequence $\{u_{n}\}_{n=1}^{\infty}\subset H^{1}_{r}(\mathbb{R}^{3})$ such that
@@
It follows from [13, Lemma 1] that
@@
Note that $\phi_{u_{n}}\rightarrow\phi_{u}$ strongly in $L^{6}(\mathbb{R}^{3})$ and $u_{n}\rightarrow u$ strongly in $L^{12/5}{(\mathbb{R}^{3})}$. Then
@@
By Lemma 1, 2 and the famous mountain pass theorem, there is a $u_{\epsilon}\in H^{1}_{r}(\mathbb{R}^{3})$ such that $I_{\epsilon}(u_{\epsilon})=c_{\epsilon}$ and $I_{\epsilon}^{\prime}(u_{\epsilon})=0$. Therefore, $u_{\epsilon}$ is the weak solution of system (2.3). Indeed, the mountain pass value is defined as follows
$c_{\epsilon}=\inf\limits_{\gamma\in\Gamma}\sup\limits_{t\in[0,1]}I_{\epsilon}(\gamma(t)),$
where
$\Gamma=\{\gamma\in C([0,1],H^{1}_{r}(\mathbb{R}^{3})):~\gamma(0)=0,\gamma(1)=\gamma_{0}(1)~\text{and}~I_{\epsilon}(\gamma_{0}(1))<0\}.$
By Lemma 1-(i) we have
(3.10) $c_{\epsilon}\geq\beta>0.$
Since $\rho$ and $\beta$ are positive constants independent of $\epsilon$. And the function $\zeta\in H^{1}_{r}(\mathbb{R}^{3})$ in Lemma 1-(ii) is also independent of $\epsilon$. Then we have the following result:
@@
There exist two positive constants $m_{1},m_{2}>0$, independent of $\epsilon\in(0,1]$, such that
$m_{1}\ \leq\ I_{\epsilon}(u_{\epsilon})=c_{\epsilon}\ \leq\ m_{2},$
where $u_{\epsilon}\in H^{1}_{r}(\mathbb{R}^{3})$ is the mountain pass critical point of $I_{\epsilon}$.
Proof.
By Lemma 1 and the Palais–Smale condition (Lemma 2), for every $\epsilon\in(0,1]$ there exists $u_{\epsilon}\in H^{1}_{r}(\mathbb{R}^{3})$ with
$I_{\epsilon}^{\prime}(u_{\epsilon})=0~\text{and}~I_{\lambda}(u_{\epsilon})=c_{\epsilon}\geq\beta>0,$
where $\beta$ is independent of $\epsilon$. Thus, we choose $m_{1}:=\beta$.
For the upper bound, consider the comparison functional (the case $\epsilon=1$)
[…]
$\leq\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x+\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\left(\epsilon^{2}\delta|u|^{2}+\epsilon^{3}C_{\delta}|u|^{3}\right)\text{d}x$
$\leq\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x+\int_{\mathbb{R}^{3}}\left(\delta|u|^{2}+C_{\delta}|u|^{3}\right)\text{d}x=J(u).$
Since $\epsilon\leq 1$, we have the point-wise comparison
$I_{\epsilon}(u)\ \leq\ J(u)\quad\text{for~all}~u\in H^{1}_{r}(\mathbb{R}^{3}).$
Choose $h\in C_{0}^{\infty}(\mathbb{R}^{3})$ with $h\neq 0$. As in the proof of Lemma 1-(ii), there exists $s\gg 1$ such that $J(sh)<0$. Fix the path $\gamma_{0}(t):=t\,sh$ on $t\in[0,1]$. Then, by definition of the mountain pass level,
$c_{\epsilon}\ =\ \inf_{\gamma\in\Gamma}\;\sup_{t\in[0,1]}I_{\epsilon}(\gamma(t))\ \leq\ \sup_{t\in[0,1]}I_{\epsilon}(\gamma_{0}(t))\leq\ \max_{t\in[0,1]}J(\gamma_{0}(t))=:m_{2}.$
The map $t\mapsto J(\gamma_{0}(t))$ is continuous on the compact interval $[0,1]$; hence the supremum is a finite maximum, and $m_{2}$ is independent of $\epsilon$. Therefore $I_{\epsilon}(u_{\epsilon})=c_{\epsilon}\leq m_{2}$, which, together with the lower bound, proves the lemma. ∎
The following lemma enables us to remove the truncation (2.1).
Lemma 4.
There exists $\epsilon^{\prime}<\epsilon_{\ast}$ such that for $0<\epsilon<\epsilon^{\prime}$, let $u_{\epsilon}$ be the weak solution of system (2.3) obtained in Lemma 2. Then there exists a positive constant $m_{0}$ independent of $\epsilon$ such that
@@
For any $0<\epsilon<\epsilon_{*}$, let $u_{\epsilon}$ satisfy $I_{\epsilon}(u_{\epsilon})=c_{\epsilon}$ and $I_{\epsilon}^{\prime}(u_{\epsilon})=0$ where $c_{\epsilon}\leq m_{2}$. Using the similar method in Lemma 2, it implies
@@
It is easy to say that $\{u_{\epsilon}\}$ is a bounded sequence. Therefore, there exists $M_{0}>0$ independent of $\epsilon$ such that for all $\epsilon\in(0,\epsilon_{*})$, it holds
@@
(3.12) $\int_{\mathbb{R}^{3}}\left(\nabla u_{\epsilon}\nabla\varphi+V(x)u_{\epsilon}\varphi\right)\text{d}x=\int_{\mathbb{R}^{3}}\phi_{u_{\epsilon}}u_{\epsilon}\varphi\text{d}x-\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\widetilde{F}^{\prime}\left(\epsilon u_{\epsilon}\right)\varphi\text{d}x.$
@@
$\quad\left|\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x-\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\widetilde{F}^{\prime}\left(\epsilon u\right)u\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x\right|$
$\leq\left(\int_{\mathbb{R}^{3}}\phi_{u}^{6}\text{d}x\right)^{\frac{1}{6}}\left(\int_{\mathbb{R}^{3}}\left|u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\right|^{\frac{6}{5}}\text{d}x\right)^{\frac{5}{6}}+\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\left|\widetilde{F}^{\prime}\left(\epsilon u\right)u\right|\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x$
@@
$\leq CM^{2}_{0}\left(\int_{\mathbb{R}^{3}}\left|u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\right|^{\frac{6}{5}}\text{d}x\right)^{\frac{5}{6}}+\delta\int_{\mathbb{R}^{3}}u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x+C_{\delta}\int_{\mathbb{R}^{3}}\epsilon\left|u\right|^{3}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x.$
Then take $\delta=\frac{V_{0}}{2}$ and $\epsilon’=\min\{C_{\frac{V_{0}}{2}}^{-1},\epsilon_{\ast}\}$ for $0<\epsilon\leq\epsilon$’, one has
$\quad\left|\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x-\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\widetilde{F}^{\prime}\left(\epsilon u\right)u\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x\right|$
$\leq C_{1}\left(\int_{\mathbb{R}^{3}}\left|w_{k}\right|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}+\frac{V_{0}}{2}\int_{\mathbb{R}^{3}}w_{k}^{2}\text{d}x+\epsilon C_{\frac{V_{0}}{2}}\int_{\mathbb{R}^{3}}w_{k}^{2}\left|u\right|\text{d}x$
$\leq C_{1}\left(\int_{\mathbb{R}^{3}}\left|w_{k}\right|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}+\epsilon C_{\frac{V_{0}}{2}}\left(\int_{\mathbb{R}^{3}}u^{6}\text{d}x\right)^{\frac{1}{6}}\left(\int_{\mathbb{R}^{3}}|w_{k}|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}+\frac{V_{0}}{2}\int_{\mathbb{R}^{3}}w_{k}^{2}\text{d}x$
@@
The proof of Theorem 1. According to Lemmas 1 and 2, for $0<\epsilon<\epsilon^{\prime}$, there exists a nontrivial weak solution $u$ satisfies
@@
Furthermore, take $\epsilon^{*}=\min\{\frac{1}{m_{0}^{2}},\epsilon^{\prime}\}$. For $0<\epsilon<\epsilon^{*}$, it follows from (3.11) that
@@
Hence, for $0<\epsilon<\epsilon^{*}$, $u$ is also the solution of system (1.2). Consequently, $\epsilon u$ is the solution of system (1.1) and $\lambda^{*}=\frac{1}{(\epsilon^{*})^{2}}$. Thus Theorem 1 has been proved.
@@ 4. Schrödinger-Poisson system with the sign-changing potential
The purpose of this section is to obtain a nontrivial solutions of system (1.1) with the sign-changing potential.
Before proving our results, we shall introduce the suitable function space where the critical points of the modified functional $I_{\epsilon,R}$ belong to. Define the working space as follows:
@@
In this case, we consider the following modified functional. For fixed $R>1$, we define a truncation function
$\eta_{R}\left(t\right)=\left\{\begin{matrix}\begin{aligned} &1,~\text{if}~\left|t\right|\leq R-1,\\ &0,~\text{if}~\left|t\right|\geq R,\end{aligned}\end{matrix}\right$.
where $\eta_{R}\left(t\right)\in C_{0}^{\infty}(\mathbb{R},\left[0,1\right])~~\text{and}~\left|\eta_{R}^{\prime}\left(t\right)\right|\leq 2$. For $\epsilon=\frac{1}{\sqrt{-\lambda}}$, we define the modified functional $I_{\epsilon,R}\left(u\right)$ as follows:
$I_{\epsilon,R}\left(u\right)=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u\right|\right)u\right)\text{d}x.$
It is easy to check that $I_{\epsilon,R}\in C^{1}\left(H_{V}^{1}\left(\mathbb{R}^{3}\right),\mathbb{R}\right)$. And for any $u,\varphi\in H_{V}^{1}\left(\mathbb{R}^{3}\right)$, we have
$\left<I^{\prime}_{\epsilon,R}\left(u\right),\varphi\right$>= $\int_{\mathbb{R}^{3}}\left(\nabla u\nabla\varphi+V(x)u\varphi\right)\text{d}x-\int_{\mathbb{R}^{3}}\phi_{u}u\varphi\text{d}x$
$-\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}(\nabla u\nabla\varphi+\tilde{V}(x)u\varphi)\text{d}x\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u\right|\right)u\right)\text{d}x$
$-\frac{1}{\epsilon}\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}\left(f\left(\epsilon\eta_{R}\left(\left|u\right|\right)u\right)\left(\eta_{R}\left(\left|u\right|\right)\varphi+\eta_{R}^{\prime}\left(\left|u\right|\right)|u|\varphi\right)\right)\text{d}x.$
@@ 4.1. Cerami condition
In order to prove Theorem 2, we firstly need to show that the functional $I_{\epsilon,R}$ satisfies the Cerami conditions. To this end, we rewrite $I_{\epsilon,R}$ in the following form
$I_{\epsilon,R}(u)$ $=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+\tilde{V}(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x$
$\indent-\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u\right|\right)u\right)\text{d}x-\frac{m}{2}\int_{\mathbb{R}^{3}}u^{2}\text{d}x.$
We supplement the definition of the following functional
$I_{0}(u)=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x.$
According to ${\bf(f)}$, we can derive the following conclusion.
@@
For any $R>1$, there exists a constant $C_{R}>0$ independent of $\epsilon$, such that for all $\left(\epsilon,u\right)\in\left[0,1\right]\times H_{V}^{1}\left(\mathbb{R}^{3}\right)$, one has
$\left|I_{\epsilon,R}\left(u\right)-I_{0}\left(u\right)\right|\leq\frac{1}{R}+C_{R}\epsilon^{p-2}$
and
$\left|\langle I_{\epsilon,R}^{\prime}\left(u\right),u\rangle-\langle I_{0}^{\prime}\left(u\right),u\rangle\right|\leq\frac{1}{R}+C_{R}\epsilon^{p-2}.$
@@
By ${\bf(f)}$, for any $\delta>0$ and $R>1$, there exists a constant $C_{\delta,R}>0$ such that $|F\left(t\right)|\leq\delta t^{2}+C_{\delta,R}\left|t\right|^{p}$ and $\left|f\left(t\right)\right|\leq\delta\left|t\right|+C_{\delta,R}\left|t\right|^{p-1}$ for $\left|t\right|\leq R$ and some $p\in(2,6)$. Then for any $u\in H_{V}^{1}\left(\mathbb{R}^{3}\right)$, consider
$\left|I_{\epsilon,R}\left(u\right)-I_{0}\left(u\right)\right|$ $=\left|\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u\right|\right)u\right)\text{d}x\right|$
$\leq\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}\left(\delta\epsilon^{2}\eta_{R}^{2}\left(\left|u\right|\right)u^{2}+C_{\delta,R}\epsilon^{p}\eta_{R}^{p}\left(\left|u\right|\right)\left|u\right|^{p}\right)\text{d}x$
$\leq\delta\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}u^{2}\text{d}x+C_{\delta,R}\epsilon^{p-2}\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}\left|u\right|^{p}\text{d}x$
$\leq\delta R+C_{\delta,R}\epsilon^{p-2}.$
Let $\delta=\frac{1}{R^{2}}$, one has
$\left|I_{\epsilon,R}\left(u\right)-I_{0}\left(u\right)\right|\leq\frac{1}{R}+C_{R}\epsilon^{p-2}.$
For the second estimate in the lemma, we can obtain it by the same method. Thus the proof is completed. ∎
Lemma 6.
Let $\{u_{n}\}_{n=1}^{\infty}\subset H^{1}_{V}(\mathbb{R}^{3})$ be a Cerami sequence for the functional $I_{\epsilon,R}$. For any $R>1$ fixed, there exists a constant $\epsilon(R)>0$ such that if $\epsilon\in(0,\epsilon(R))$, then $u_{n}\rightarrow u\text{ in }H^{1}_{V}(\mathbb{R}^{3})$ with $u\in H^{1}_{V}(\mathbb{R}^{3})$.
Proof.
Let $\{u_{n}\}$ be a Cerami sequence of $I_{\epsilon,R}$, that is, for some $c\in\mathbb{R}$,
$I_{\epsilon,R}(u_{n})\to c~\text{and}~(1+\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})})I_{\epsilon,R}^{\prime}(u_{n})\to 0\ ~\text{as}~n\to+\infty.$